OpenStudy (anonymous):

related rates problem need help stat! Please! A cylindrical tank with radius 5m is being filled with water at a rate of 3m/min. How fast is the height of the water increasing.

OpenStudy (amistre64):

what is the formula for the volume of a cylindar?

OpenStudy (anonymous):

I believe the formula is pi r^2h

OpenStudy (anonymous):

Thank you for helping me out! I appreciate it. how do i go about solving this type \of problem.

OpenStudy (amistre64):

good, then lets take an implicit derivative wrt.time of:\[V=\pi r^2 h\]

OpenStudy (amistre64):

\[V=\pi r^2 h\] \[V'=2r~\pi r' h+\pi r^2 h'\] since r is a constant at 5, r' is 0 \[V'=\pi r^2 h'\] they give us the change in volume to be 3, and r=5 \[3=\pi 5^2 h'\] solve for h'

OpenStudy (anonymous):

Thank you so much sorry i couldn't reply back my computer froze.But i see what you did there but to solve for h what would i do?

OpenStudy (anonymous):

Sorry about that i just reread the problem and it reads 3m^3 not 3m. sorry for the typo.

OpenStudy (amistre64):

thats still a volume rate ... so 3 is fine

OpenStudy (amistre64):

if you want to include it then 3m^3 = (5m)^2 pi h' 3m^3 --------- = h' (5m)^2 pi 3 m --------- = h' 25 pi

OpenStudy (anonymous):

okay that's great I thought that that would effect the result of the problem.

OpenStudy (amistre64):

nah, but its prolly a good idea to include it to make sure the dimensions pan out :)

OpenStudy (anonymous):

oh okay so when solving for a problem like this could you just leave it in a fraction or should you simplify the answer?

OpenStudy (amistre64):

depends on who is grading it really ...

OpenStudy (amistre64):

i usually give exact and then approximate it as a decimal:\[this\approx n.xxx\]

OpenStudy (anonymous):

well in that case my professor rarely stresses simplifying so i think that leaving it in a fraction is fine. Thank you for your help amistre64. have a great day!:)

OpenStudy (amistre64):

youre welcome, and good luck :) the strategy i use is to find a formula that can be derived; take implicits and start filling in the parts that can be knowned